OpenStudy (anonymous):

A launched rocket has an altitude, in meters, given by the polynomial h+vt-4.9^2, where h is the height, in meters, from which the launch occurs, at velocity v meters per second, and t is the number of seconds for which the rocket is airborne. If a rocket is launched from the top of a tower 140 meters high with an initial upward speed of 60 meters per second, what will its height be after 3 seconds.

OpenStudy (anonymous):

oh yah physics. i like doing this :D altitude = A A = h + vt - 4.9t^2 A = 140 + 60m/s(3) - 4.9(3)^2, use units when writing solution

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

no prob. keep in mind solution should be in meters. I'm just happy its not a angular velocity prob O.o. those are annoying. also this might have been more appropriate in the physics section, but whatever?

OpenStudy (anonymous):

im still lost do i just inplug every number or what do i do